OpenStudy (anonymous):

Please check my answer(: A right triangle has legs with lengths 7 cm and 24 cm. What is the length of the hypotenuse? ___cm (23) ? thanks btw <3 **no im not gay yes im a guy** -.-

OpenStudy (anonymous):

please help meh D:

OpenStudy (acxbox22):

what is your asnwer u need us to check?

OpenStudy (anonymous):

mmy answer is 23

OpenStudy (anonymous):

How'd you get 23? show us your work

OpenStudy (anonymous):

and yep

OpenStudy (anonymous):

ugh jeez, 24^2 - 7^2 is 527 and the square root for that is 23 (in the tenth place..rounded)

OpenStudy (anonymous):

\[A^2+B^2 = C^2 ...............7^{2} + 23^2 = C^2\]

OpenStudy (anonymous):

sooo.....

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

:) Come on, i laid it out for you. 49 + 576 = c^2

OpenStudy (anonymous):

The problem was that you messed up the formula. It's A2+B2=C2. Not A2-B2=C2.

OpenStudy (anonymous):

331776 e.e

OpenStudy (anonymous):

Ah...

OpenStudy (anonymous):

idk if thats right.

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

\[A ^{2}+B^{2}=C^{2} \rightarrow 7^{2} + 24^{2} = C^{2} \rightarrow 49 + 576 = C^{2} \rightarrow 625 = C^{2} \] ..... Now solve for C

OpenStudy (anonymous):

um o.o dude im lostt.

OpenStudy (anonymous):

im just gonna go with 23 o.o who cares, right XD

OpenStudy (anonymous):

just kidding XD

OpenStudy (anonymous):

Okay. First of all, i typed 24 instead of 23. My bad. The way to solve this is to use the Pythagorean Theorem. The therom says that \[A^2 + B^2 = C^2\] .Take both legs given to the second power. This gives you \[7^2 + 24^2 = C^2\]. Now once you have the numbers plugged in, you're left with \[49 + 576 = C^2\]. From here you got: \[625 = C^2\] Finally, you solve for C. Which you take the square of to counter the power. \[\sqrt{625} = \sqrt{C^2} \] Which turns into: \[25 = C\]. Answer is 25.

OpenStudy (anonymous):

oh my god dude, thanks I luhh you (no homo) but i luh you

OpenStudy (anonymous):

Does that make more sense?

OpenStudy (anonymous):

Yea no problem, just make sure you're using the right formulas. You were almost there but you stumbled at the end.

OpenStudy (anonymous):

yeppp.